For a binomial distribution, the probability of success on any trial is 35 and 240 trials were carried out. What is the standard deviation of the distribution of number of points?


1
Expert's answer
2021-02-24T07:29:12-0500

We have that

n = 240

p = 0.35

Standard deviation is calculated by the formula:


"\sigma = \sqrt{npq}"

where q = 1 – p = 1 – 0.35 = 0.65

Thus

"\sigma = \sqrt{240\cdot 0.35 \cdot 0.65} = 7.39"


Answer: 7.39